# Bayesian Hierarchical Probabilistic Forecasting of Intraday Electricity Prices

### v1

@@ Abstract
We present a first study of Bayesian forecasting of electricity prices traded on the German continuous intraday market which fully incorporates parameter uncertainty. Our target variable is the IDFull price index, forecasts are given in terms of posterior predictive distributions. For validation we use the exceedingly volatile electricity prices of 2022, which have hardly been the subject of forecasting studies before. As a benchmark model, we use all available intraday transactions at the time of forecast creation to compute a current value for the IDFull. According to the weak-form efficiency hypothesis, it would not be possible to significantly improve this benchmark built from last price information. We do, however, observe statistically significant improvement in terms of both point measures and probability scores. Finally, we challenge the declared gold standard of using LASSO for feature selection in electricity price forecasting by presenting strong statistical evidence that Orthogonal Matching Pursuit (OMP) leads to better forecasting performance.
@@ 1 Introduction and Motivation
In May 2023 the monthly share of renewable electricity generation in the European Union reached 49.3% (energy-charts.info), a new high and fast approaching the milestone of having a majority of electrical energy produced by renewables in the EU. This milestone has already been reached March 2019 in Germany’s energy generation, now approaching a yearly share of 60% renewables. But the penetration of Renewable Energy Sources (RES) necessary for achieving the vision of sustainable energy production comes with tough challenges, as, opposed to other sources of energy, wind and solar energy production cannot be planned in advance, and mechanisms need to be in place to deal with the intermittency of RES in order to keep our energy systems stable.
@@
Point forecasts of electricity prices are typically based on regularised regression models (Kremer et al., 2020; Kath and Ziel, 2018), in particular autoregressive models (Hu et al., 2021; Lucic and Xydis, 2023), as well as artificial neural networks (Oksuz and Ugurlu, 2019; Lehna et al., 2022). However, in view of the strong volatility in the CID market, in particular in Germany with their high solar and wind energy components in the energy mix (c.f. Figure 1), most of the recent works on CID EPF agree upon the importance of probabilistic forecasting (Maciejowska et al., 2023b). Employed probabilistic forecasting methods include generalised additive models for location, scale and shape (GAMLSS) (Abramova and Bunn, 2020; Narajewski and Ziel, 2020b; Narajewski, 2022; Hirsch and Ziel, 2022, 2024), distributional neural networks (DDNNs) (Marcjasz et al., 2023; Narajewski, 2022; Barunik and Hanus, 2023), quantile regression averaging (QRA) (Marcjasz et al., 2023; Maciejowska et al., 2023a; Andrade et al., 2017; Uniejewski and Weron, 2021; Serafin et al., 2019), and other specialised techniques (Cramer et al., 2023; Grothe et al., 2023). However, all of these approaches use point estimates of distribution parameters (or quantiles) to build a probabilistic forecast model, and as such do not account fully for the uncertainty of parameter estimation. With our work, we fill this gap by promoting a Bayesian forecast model tested on the German CID market for the recent period of 2021 - 2022 featuring particularly volatile electricity prices.
@@
A crucial step of EPF is the choice of regressors. Typical regressors found in most works are price and volume information of the electricity markets, external variables like forecasted power generation of various energy sources and load forecasts, and dummy time and date variables to capture seasonal trends. But also carbon emission allowances (Pape et al., 2016; Marcjasz et al., 2023; Maciejowska et al., 2020, 2023a), unavailability of power generation (Pape et al., 2016; Hirsch and Ziel, 2022; Lima et al., 2023), cross-border flow of energy (Pape et al., 2016; Andrade et al., 2017), trade closing times (Hirsch and Ziel, 2022, 2024), market elasticity (Kremer et al., 2021; Hirsch and Ziel, 2022, 2024), and grid frequency (Scholz et al., 2021) have been utilised. The market elasticity is approximated by the slopes of supply and demand curves through two-sided auctions in the day-ahead market (Kulakov and Ziel, 2019). Typically, also autoregressive components are considered as extra regressors (Maciejowska et al., 2019, 2021; Janke and Steinke, 2019; Kath, 2019; Maciejowska et al., 2020; Uniejewski and Weron, 2021; Uniejewski et al., 2019). These ARX-type models typically use lags of one day or one hour. Often, a lag of two days and, to account for weekly seasonality, seven days is also considered, but these lags have been reported as less relevant (Pape et al., 2016; Shinde et al., 2021).
In view of this large set of regressors, possibly with strong co-linearities, selection of regressors or features is crucial. In the EPF, the Least Absolute Shrinkage and Selection Operator (LASSO) (Tibshirani, 1996) is very popular and has even been declared the gold standard of EPF (Maciejowska et al., 2023b). In this work, referring to the known instabilities the LASSO has when facing strong co-linearities (Su et al., 2017), we challenge this standard and instead promote Orthogonal Matching Pursuit (OMP) (Tropp and Gilbert, 2007; Rubinstein et al., 2008; Pedregosa et al., 2011) as a feature selection technique for the case of large sets of regressors.
@@
    To the best of our knowledge, we present the first complete Bayesian treatment of CID EPF fully incorporating uncertainties of model parameters.
@@
    Apart from the recent work Hirsch and Ziel (2024), we are not aware of any other work concerned with the CID prices for the exceedingly volatile years 2021 and 2022.
@@
  4. 4.
    We add to the discussion of the proposed weak-form efficiency of CID markets.
  5. 5.
    We share a detailed description of CID indices and statistics calculations which has become more intricate since 2021, and, as far as we are aware, has not yet been published to this detail.
  6. 6.
    We leverage state-of-the-art probabilistic methods in python, i.e. TensorFlow Probability (TFP) (Dillon et al., 2017), and the acceleration and parallelisation environment JAX (Bradbury et al., 2018).
@@ 2 Electricity markets and market data
Understanding the characteristics of the European electricity market is crucial for establishing successful forecasting models. In this section, we therefore summarise the market details important for our data model. More details of the (German) short-term electricity market can be found in Viehmann (2017) and other references given.
@@
The CID market is one of many platforms to trade electricity that resulted from the liberalisation of the European energy sector since 1996 (europarl.europa.eu). The most liquid market is the day-ahead (DA) market conducting a uniform price auction to settle clearance prices of electricity for the next day. Markets like the CID and DA markets in the European trading zone are operated by Nominated Electricity Market Operators (NEMOs), of which EPEX is one of 17 (nemo-committee.eu). However, due to the large share of volumes traded, EPEX is generally considered as the reference point for electricity prices in Germany and other countries (Viehmann, 2017). In addition to NEMOs managing the trading platforms, various Transmission System Operators (TSOs) share the responsibility for the transmission of electrical power and for providing market participants access to the grid (entsoe.eu).
@@ 2.4 Live intraday values
We therefore use the ExecutionTime information of transactions to take the perspective of a market participant trading power on the CID market at a thought point of time, and filter out any transactions that have taken place after this time. Thus, we simulate the forecast creation time $\tau$ at an arbitrary point of time, and forecast the IDFull of all following hours. To indicate preliminary values for CID price indices and statistics, we will use the prefix “live”, e.g. live IDFull.
@@ 3 The data model
Regressor
Definition
Source
Day-ahead market information EPEX Spot
$P_{\mathrm{da}}$
Market-clearance price for SDAC countries (DA price, spot price).
$P_{\mathrm{da}}^{\mathrm{che}}$
Market-clearance price for Switzerland (non-SDAC).
$V_{\mathrm{da}}$
Market-clearance volume for EPEX Germany.
$P_{\mathrm{da}}^{15,1}$, …, $P_{\mathrm{da}}^{15,4}$
@@
The four quarter-hourly SDAC market clearance prices of delivery hour $h$.
$s_{15}(P_{\mathrm{da}})$
@@
Average slope quarter-hourly SDAC market clearance prices, i.e. $s_{15}(P_{\mathrm{da}})=\frac{1}{3}\Big{(}P_{\mathrm{da}}^{15,4}-P_{\mathrm{da}}^{15,1}\Big{)}$
$\eta_{\mathrm{da}}(\Delta p)$
@@
Merit-order slopes at $P_{\mathrm{da}}$ calculated for three different finite differences $\Delta p=500,1000,2000$ as measure of elasticity.
@@
Statistical day-ahead market information from other hours (EPEX moderated) ENTSO-E
$P_{\mathrm{da}}(h_{1},h_{2},...)$
@@
Average DA price for selected delivery hours $h_{1},h_{2},...$ representing middle-night, early morning, morning, late morning, high noon, early afternoon, afternoon, evening, night, baseload, off-peak, rush hour, sunpeak, peakload, maximum, minimum, volume-weighted average.
Continuous intraday market information EPEX Spot
$P_{\mathrm{id1}}$, $P_{\mathrm{id3}}$, $P_{\mathrm{idfull}}$
Live CID price indices ID1, ID3 and IDFull.
$P_{\mathrm{high}}$, $P_{\mathrm{low}}$, $P_{\mathrm{last}}$
Live CID price statistics High, Low, Last.
$P_{\mathrm{deviat}}$
Live volume-weighted average deviation from mean price, i.e.
$P_{\mathrm{deviat}}=\frac{1}{\sum_{j\in I}v_{j}}\sum_{i\in I}v_{i}\,\Big{[}p_{i}-\frac{1}{|I|}\sum_{j\in I}p_{j}\Big{]},$
@@
where $\{(p_{i},v_{i})\,|\,i\in I\}$ are price and volume tuples of transactions executed before forecast creation.
$V_{\mathrm{buy}}$, $V_{\mathrm{sell}}$
@@
Live sums of volumes purchased and sold on the CID market.
$\mathrm{eod}_{d-1}(x)$
@@
Final values of indices and statistics for day $d-1$, where $x=P_{\mathrm{id1}},\,P_{\mathrm{id3}},\,y,\,P_{\mathrm{high}},\,P_{\mathrm{low}},\,P_{\mathrm{last}},\,P_{\mathrm{deviat}},\,V_{\mathrm{buy}},\,V_{\mathrm{sell}}$.
$\eta_{\mathrm{cid}}(\Delta p)$
@@
Merit-order slopes at $P_{\mathrm{id1}}$ calculated for three different finite differences $\Delta p=500,1000,2000$ as measure of elasticity.
@@
Table 1: Base market regressors and their notation used in the model. If not stated differently, all variables pertain delivery day $d$ for a fixed hour $h$ in Germany. Forecasts and live values depend on availability at forecast creation time $\tau$. External regressors are listed in Table 2, additional features are calculated from these regressors and are listed in Table 3.
Regressor
Definition
Source
Energy consumption and production forecasts for Germany ENTSO-E
$E_{\mathrm{cons}}$
@@
Energy consumption calculated as the sum of day-ahead forecasts of load for the four 15 min periods of delivery hour $h$.
$E_{\mathrm{sol}},E_{\mathrm{won}},E_{\mathrm{woff}}$
@@
Solar, wind onshore and wind offshore energy production calculated as the sum of the four 15 min power forecasts of delivery hour $h$, where intraday forecasts replace day-ahead forecasts after 8:00 am on day $d$.
$E_{\mathrm{tot}}$
@@
Total energy production calculated as the sum of the four 15 min day-ahead power forecasts of delivery hour $h$.
@@
Energy consumption and production forecasts for other countries ENTSO-E
@@
Date and time dummy variables python datetime
@@
$T_{\mathrm{deliv}}=h-\tau$
Time to delivery in floating point hours.
$T_{\mathrm{public}}$
Population-weighted number of states in work holiday.
arbeitstage.org
$T_{\mathrm{school}}$
Population-weighted number of states in school holiday.
ferienwiki.de
@@
$M_{\mathrm{gas}}$
S&P GSCI Natural Gas index.
S&P Global
$M_{\mathrm{gasoil}}$
S&P GSCI Gasoil index.
S&P Global
@@
$S_{\mathrm{tnoon}}$
Time to noon in floating point hours.
@@
$S_{\mathrm{dl}}$
Lengths of days as time from sunrise to sunset.
$S_{\mathrm{temp}}$
Yearly season indicator constructed from 15-year hourly temperature averages across weather stations in Germany.
meteostat
Table 2: External base regressors and their notation, continuation of Table 1.
@@ 3.2 External regressors
Standard external variables are power generation and load forecasts publicly available from the transparency platform of European Network of Transmission System Operators for Electricity (ENTSO-E) (entsoe.eu). These include day-ahead forecasts, i.e. created 18:00 pm on day $d-1$, of power generation from a multitude of different energy sources, as well as load forecasts. In addition, intraday forecasts, i.e. created 8:00 am on day $d$, are published for offshore and onshore wind power and solar power. Some forecasts pertain 15 min periods, in these cases we aggregate all power forecasts to hourly energy productions, as well as to average slopes of 15 min power forecasts when available. Depending on what is available at forecast creation, we pick day-ahead or intraday forecasts.
Overall, we have hourly consumption $E_{\mathrm{cons}}$ derived from load forecasts, and hourly renewable energy productions comprising solar, $E_{\mathrm{sol}}$, onshore wind $E_{\mathrm{won}}$, and offshore wind $E_{\mathrm{woff}}$, as well as the total hourly energy production $E_{\mathrm{tot}}$ from power forecasts. Using the notation $s_{15}(x)$ for the 15 min slope of the corresponding variable $x$, we also consider $s_{15}(E_{\mathrm{cons}})$, $s_{15}(E_{\mathrm{sol}})$, $s_{15}(E_{\mathrm{won}})$ and $s_{15}(E_{\mathrm{woff}})$. To also take cross-border effects on price formation into account, we include power generation forecasts of Slovenia, Switzerland, Italy, Hungary and Czechia.
@@
An important aspect for EPF, especially for the CID market in recent years, are market states varying with time. Events strongly affecting the markets in 2021 and 2022 obviously arose in the course of the COVID-19 pandemic and the Russian invasion of Ukraine, but also smaller events like interest policies and legislations related to energy can play a major role. These changing market states impede training of forecast model with historical data as stationary dynamics is not given. In particular the years 2021 and 2022 show extreme fluctuations due to major changes in market states, c.f. Figure 1. Ideally, custom variables capturing all dimensions of market states may be incorporated as regressors, but in view of the difficulty of this task, we take the S&P GSCI Natural Gas and S&P GSCI Gasoil indices $M_{\mathrm{gas}}$ and $M_{\mathrm{gasoil}}$ as approximations for market states. These index values are of daily resolution and exclude weekends and holidays, we therefore always take the most recent available value with respect to delivery day and forecast creation.

### v2

@@ Abstract
We present a first study of Bayesian forecasting of electricity prices traded on the German continuous intraday market which fully incorporates parameter uncertainty. A particularly large set of endogenous and exogenous covariables is used, handled through feature selection with Orthogonal Matching Pursuit (OMP) and regularising priors. Our target variable is the IDFull price index, forecasts are given in terms of posterior predictive distributions. For validation we use the exceedingly volatile electricity prices of 2022, which have hardly been the subject of forecasting studies before. As a benchmark model, we use all available intraday transactions at the time of forecast creation to compute a current value for the IDFull. According to the weak-form efficiency hypothesis, it would not be possible to significantly improve this benchmark built from last price information. We do, however, observe statistically significant improvement in terms of both point measures and probability scores. Finally, we challenge the declared gold standard of using LASSO for feature selection in electricity price forecasting by presenting strong statistical evidence that OMP leads to better forecasting performance.
@@ 1 Introduction and Motivation
In April 2024 the monthly share of renewable electricity generation in the European Union reached 52.3% (energy-charts.info), a new high and for the first time surpassing the milestone of having a majority of electrical energy produced by renewables in the EU. This milestone has already been reached March 2019 in Germany’s energy generation, now approaching a yearly share of 60% renewables. But the penetration of Renewable Energy Sources (RES) necessary for achieving the vision of sustainable energy production comes with tough challenges, as, opposed to other sources of energy, wind and solar energy production cannot be planned in advance, and mechanisms need to be in place to deal with the intermittency of RES in order to keep our energy systems stable.
@@
Point forecasts of electricity prices are typically based on regularised regression models (Kremer et al., 2020; Kath and Ziel, 2018), in particular autoregressive models (Hu et al., 2021; Lucic and Xydis, 2023), as well as artificial neural networks (Oksuz and Ugurlu, 2019; Lehna et al., 2022), or, more in the context of electricity demand forecast, functional data approaches Vilar et al. (2012); Shah et al. (2022); Varelas et al. (2024).
However, in view of the strong volatility in the CID market, in particular in Germany with their high solar and wind energy components in the energy mix (cf. Figure 1), most of the recent works on CID EPF agree upon the importance of probabilistic forecasting (Maciejowska et al., 2023b). Employed probabilistic forecasting methods include generalised additive models for location, scale and shape (GAMLSS) (Abramova and Bunn, 2020; Narajewski and Ziel, 2020b; Narajewski, 2022; Hirsch and Ziel, 2022, 2024), distributional neural networks (DDNNs) (Marcjasz et al., 2023; Narajewski, 2022; Barunik and Hanus, 2023), quantile regression averaging (QRA) (Marcjasz et al., 2023; Maciejowska et al., 2023a; Andrade et al., 2017; Cabrera and Schulz, 2017; Uniejewski and Weron, 2021; Serafin et al., 2019), and other specialised techniques (Cramer et al., 2023; Grothe et al., 2023). However, all of these approaches use point estimates of distribution parameters (or quantiles) to build a probabilistic forecast model, and as such do not account fully for the uncertainty of parameter estimation. With our work, we fill this gap by promoting a Bayesian forecast model tested on the German CID market for the recent period of 2021 - 2022 featuring particularly volatile electricity prices.
@@
A crucial step of EPF is the choice of regressors. Typical regressors found in most works are price and volume information of the electricity markets, external variables like forecasted power generation of various energy sources and load forecasts, and dummy time and date variables to capture seasonal trends. But also carbon emission allowances (Pape et al., 2016; Marcjasz et al., 2023; Maciejowska et al., 2020, 2023a), unavailability of power generation (Pape et al., 2016; Hirsch and Ziel, 2022; Lima et al., 2023), cross-border flow of energy (Pape et al., 2016; Andrade et al., 2017), trade closing times (Hirsch and Ziel, 2022, 2024), market elasticity (Kremer et al., 2021; Hirsch and Ziel, 2022, 2024), and grid frequency (Scholz et al., 2021) have been utilised. The market elasticity is approximated by the slopes of supply and demand curves through two-sided auctions in the day-ahead market (Kulakov and Ziel, 2019). Typically, also autoregressive components are considered as additional regressors (Maciejowska et al., 2019, 2021; Janke and Steinke, 2019; Kath, 2019; Maciejowska et al., 2020; Uniejewski and Weron, 2021; Uniejewski et al., 2019). These ARX-type models typically use lags of one day or one hour. Often, a lag of two days and, to account for weekly seasonality, seven days is also considered, but these lags have been reported as less relevant (Pape et al., 2016; Shinde et al., 2021).
In view of this large set of regressors, possibly with strong collinearities, selection of regressors or features is crucial. The Least Absolute Shrinkage and Selection Operator (LASSO) (Tibshirani, 1996) is a popular choice and has been declared the gold standard of EPF (Maciejowska et al., 2023b). In this work, referring to the known instabilities the LASSO has when facing strong collinearities (Su et al., 2017), we challenge this standard and instead promote Orthogonal Matching Pursuit (OMP) (Tropp and Gilbert, 2007; Rubinstein et al., 2008; Pedregosa et al., 2011) as a feature selection technique for the case of large sets of regressors with strong correlations.
@@
    To the best of our knowledge, we present the first complete Bayesian treatment of CID EPF, fully incorporating uncertainties of model parameters. The study period is the exceedingly volatile years 2021 and 2022, which, apart from the recent work by Hirsch and Ziel (2024), has not been the subject of CID EPF before as far as we are aware.
@@
    We add to the discussion of the proposed weak-form efficiency of CID markets and share a detailed description of CID indices and statistics calculations which have become more intricate since 2021, and, as far as we are aware, have not yet been published in this detail.
@@ 2 Electricity markets and market data
Understanding the characteristics of the European electricity market is crucial for establishing successful forecasting models. In this section, we therefore summarise the market details important for our data model. More details of the (German) short-term electricity market can be found in Viehmann (2017) and the references given below.
@@
The CID market resulted from the liberalisation of the European energy sector since 1996 (europarl.europa.eu) and is one of many platforms to trade electricity. The most liquid market is the day-ahead (DA) market conducting a uniform price auction to settle clearance prices of electricity for the next day. Markets like the CID and DA markets in the European trading zone are operated by Nominated Electricity Market Operators (NEMOs), of which EPEX is one of 17 (nemo-committee.eu). However, due to the large share of volumes traded, EPEX is generally considered as the reference point for electricity prices in Germany and other countries (Viehmann, 2017). In addition to NEMOs managing the trading platforms, various Transmission System Operators (TSOs) share the responsibility for the transmission of electrical power and for providing market participants access to the grid (entsoe.eu).
@@ 2.4 Live intraday values
We therefore use the ExecutionTime information of transactions to take the perspective of a market participant trading power on the CID market at a thought point of time, and filter out any transactions that have taken place after this time. Thus, we emulate a forecast creation time $\tau$ and forecast the IDFull of all following hours. To indicate these preliminary values for CID price indices and statistics, we will use the prefix “live”, e.g. live IDFull.
@@ 3 The data model
Regressor Definition Source
Day-ahead market information I EPEX Spot
$P_{\mathrm{da}}$ Market-clearance price for SDAC countries (DA price, spot price).
$P_{\mathrm{da}}^{\mathrm{che}}$ Market-clearance price for Switzerland (non-SDAC).
$V_{\mathrm{da}}$ Market-clearance volume for EPEX Germany.
@@
$P_{\mathrm{da}}^{15,1}$, …, $P_{\mathrm{da}}^{15,4}$ The four quarter-hourly SDAC market clearance prices of delivery hour $h$.
@@
$s_{15}(P_{\mathrm{da}})$ Average slope quarter-hourly SDAC market clearance prices, i.e. $s_{15}(P_{\mathrm{da}})=\frac{1}{3}\Big{(}P_{\mathrm{da}}^{15,4}-P_{\mathrm{da}}^{15,1}\Big{)}$
@@
$\eta_{\mathrm{da}}(\Delta p)$ Merit-order slopes at $P_{\mathrm{da}}$ calculated for three different finite differences $\Delta p=500,1000,2000$ as measure of elasticity.
@@
Statistical day-ahead market information from other hours (EPEX moderated) I ENTSO-E
@@
$P_{\mathrm{da}}(h_{1},h_{2},...)$ Average DA price for selected delivery hours $h_{1},h_{2},...$ representing middle-night, early morning, morning, late morning, high noon, early afternoon, afternoon, evening, night, baseload, off-peak, rush hour, sunpeak, peakload, maximum, minimum, volume-weighted average.
Continuous intraday market information I EPEX Spot
$P_{\mathrm{id1}}$, $P_{\mathrm{id3}}$, $P_{\mathrm{idfull}}$ Live CID price indices ID1, ID3 and IDFull.
$P_{\mathrm{high}}$, $P_{\mathrm{low}}$, $P_{\mathrm{last}}$ Live CID price statistics High, Low, Last.
@@
$P_{\mathrm{deviat}}$ Live volume-weighted average deviation from mean price, i.e. $P_{\mathrm{deviat}}=\frac{1}{\sum_{j\in I}v_{j}}\sum_{i\in I}v_{i}\,\Big{[}p_{i}-\frac{1}{|I|}\sum_{j\in I}p_{j}\Big{]},$ where $\{(p_{i},v_{i})\,|\,i\in I\}$ are price and volume tuples of transactions executed before forecast creation.
@@
$V_{\mathrm{buy}}$, $V_{\mathrm{sell}}$ Live sums of volumes purchased and sold on the CID market.
@@
$\mathrm{eod}_{d-1}(x)$ Final values of indices and statistics for day $d-1$, where $x=P_{\mathrm{id1}},\,P_{\mathrm{id3}},\,y,\,P_{\mathrm{high}},\,P_{\mathrm{low}},\,P_{\mathrm{last}},\,P_{\mathrm{deviat}},\,V_{\mathrm{buy}},\,V_{\mathrm{sell}}$.
@@
$\eta_{\mathrm{cid}}(\Delta p)$ Merit-order slopes at $P_{\mathrm{id1}}$ calculated for three different finite differences $\Delta p=500,1000,2000$ as measure of elasticity.
@@
Table 1: Base market regressors and their notation used in the model. If not stated differently, all variables pertain delivery day $d$ for a fixed hour $h$ in Germany. Forecasts and live values depend on availability at forecast creation time $\tau$. The background of these variables is explained in Section 2. External regressors are listed in Table 2, additional features are calculated from these regressors and are listed in Table 3.
Regressor Definition Source
Energy consumption and production forecasts for Germany I ENTSO-E
@@
$E_{\mathrm{cons}}$ Energy consumption calculated as the average of day-ahead forecasts of load for the four 15 min periods of delivery hour $h$.
@@
$E_{\mathrm{sol}},E_{\mathrm{won}},E_{\mathrm{woff}}$ Solar, wind onshore and wind offshore energy production calculated as the average of the four 15 min power forecasts of delivery hour $h$, where intraday forecasts replace day-ahead forecasts after 8:00 am on day $d$.
@@
$E_{\mathrm{tot}}$ Total energy production calculated as the average of the four 15 min day-ahead power forecasts of delivery hour $h$.
@@
Energy consumption and production forecasts for other countries I ENTSO-E
@@
Date and time dummy variables I python datetime
@@
$T_{\mathrm{deliv}}=h-\tau$ Time to delivery in floating point hours.
$T_{\mathrm{public}}$ Population-weighted number of states in work holiday. arbeitstage.org
$T_{\mathrm{school}}$ Population-weighted number of states in school holiday. ferienwiki.de
@@
$M_{\mathrm{gas}}$ S&P GSCI Natural Gas index. S&P Global
$M_{\mathrm{gasoil}}$ S&P GSCI Gasoil index. S&P Global
@@
$S_{\mathrm{tnoon}}$ Time to noon in floating point hours.
@@
$S_{\mathrm{dl}}$ Lengths of days as time from sunrise to sunset.
$S_{\mathrm{temp}}$ Yearly season indicator constructed from 15-year hourly temperature averages across weather stations in Germany. meteostat
Table 2: External base regressors and their notation, continuation of Table 1. The variables listed here are explained in Section 3.2.
@@ 3.2 External regressors
Standard external variables are power generation and load forecasts publicly available from the transparency platform of the European Network of Transmission System Operators for Electricity (ENTSO-E) (entsoe.eu). These include day-ahead forecasts, i.e. created 18:00 pm on day $d-1$, of power generation from a multitude of different energy sources, as well as load forecasts. In addition, intraday forecasts, i.e. created 8:00 am on day $d$, are published for offshore and onshore wind power and solar power. Some forecasts pertain 15 min periods, in these cases we aggregate all power forecasts to hourly energy productions, as well as to average slopes of 15 min power forecasts when available. Depending on what is available at forecast creation, we pick day-ahead or intraday forecasts.
Overall, we have hourly consumption $E_{\mathrm{cons}}$ derived from load forecasts, and hourly renewable energy productions comprising solar $E_{\mathrm{sol}}$, onshore wind $E_{\mathrm{won}}$, and offshore wind $E_{\mathrm{woff}}$, as well as the total hourly energy production $E_{\mathrm{tot}}$ from power forecasts. Using the notation $s_{15}(x)$ for the 15 min slope of the corresponding variable $x$, we also consider $s_{15}(E_{\mathrm{cons}})$, $s_{15}(E_{\mathrm{sol}})$, $s_{15}(E_{\mathrm{won}})$ and $s_{15}(E_{\mathrm{woff}})$. To also take cross-border effects on price formation into account, we include power generation forecasts of Slovenia, Switzerland, Italy, Hungary and Czechia.
@@
An important aspect for EPF, especially for the CID market in recent years, are market states varying with time. Events strongly affecting the markets in 2021 and 2022 obviously arose in the course of the COVID-19 pandemic and the Russian invasion of Ukraine, but also smaller events like interest policies and legislations related to energy can play a major role. These changing market states impede training of forecast model with historical data as stationary dynamics is not given. In particular the years 2021 and 2022 show extreme fluctuations due to major changes in market states, cf. Figure 1. Ideally, custom variables capturing all dimensions of market states may be incorporated as regressors, but in view of the difficulty of this task, we take the S&P GSCI Natural Gas and S&P GSCI Gasoil indices $M_{\mathrm{gas}}$ and $M_{\mathrm{gasoil}}$ as approximations for market states. These index values are of daily resolution and exclude weekends and holidays, we therefore always take the most recent available value with respect to delivery day and forecast creation.